OpenStudy (anonymous):

Mindf**k Question- Can you guess it? Billy has 16 coins in a jar. The jar contains only nickels and dimes. If the coins are worth $1.11, how many nickels does Billy have?

OpenStudy (anonymous):

10 nickles and 6 dimes.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Lets let x=number of nickels. Since there are 16 total coins, (16-x)=number of dimes. We know the value of the coins is 1.10. So, 10 times the number of dimes plus 5 times the number of nickels equals 110 cents. Or: \[5x+10(16-x)=110\] or- \[5x+160-10x=110\] combine like terms and simplify and you find \[x=10 nickels\]

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

look again..

OpenStudy (anonymous):

You just changed the question.

OpenStudy (anonymous):

oh boo! It isnt possible! you would need a penny!

OpenStudy (anonymous):

With the value now 1.11, you'd find \[x=10.2 nickels\] which wouldn't be possible. The question is unsolvable.

OpenStudy (anonymous):

No I didnt, the question wasnt changed. And @britanny_ is getting somewhere.

OpenStudy (anonymous):

Wanna know the answer? It's dumb. My friend told me.

OpenStudy (anonymous):

when he was drunk...

OpenStudy (anonymous):

You did, in fact, change the question, the value was originally 1.10.

OpenStudy (anonymous):

i corrected a mistake i made.

OpenStudy (anonymous):

There were 17 coins. He found a penny taped to the bottom.

OpenStudy (anonymous):

Thanks for playing...

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

GG c:

OpenStudy (anonymous):

share with your friends.